# Off-Path TCP Hijacking in Wi-Fi Networks: A Packet-Size Side Channel Attack

### v4

@@ Abstract
In this paper, we unveil a fundamental side channel in Wi-Fi networks, specifically the observable frame size, which can be exploited by attackers to conduct TCP hijacking attacks. Despite the various security mechanisms (e.g., WEP and WPA2/WPA3) implemented to safeguard Wi-Fi networks, our study reveals that an off path attacker can still extract sufficient information from the frame size side channel to hijack the victim’s TCP connection. Our side channel attack is based on two significant findings: (i) response packets (e.g., ACK and RST) generated by TCP receivers vary in size, and (ii) the encrypted frames containing these response packets have consistent and distinguishable sizes. By observing the size of the victim’s encrypted frames, the attacker can detect and hijack the victim’s TCP connections. We validate the effectiveness of this side channel attack through two case studies, i.e., SSH DoS and web traffic manipulation. Precisely, our attack can terminate the victim’s SSH session in 19 seconds and inject malicious data into the victim’s web traffic within 28 seconds. Furthermore, we conduct extensive measurements to evaluate the impact of our attack on real-world Wi-Fi networks. We test 30 popular wireless routers from 9 well-known vendors, and none of these routers can protect victims from our attack. Besides, we implement our attack in 80 real-world Wi-Fi networks and successfully hijack the victim’s TCP connections in 75 (93.75%) evaluated Wi-Fi networks. We have responsibly disclosed the vulnerability to the Wi-Fi Alliance and proposed several mitigation strategies to address this issue.
1 Introduction
@@
Our attack consists of four steps. First, the attacker accesses a public Wi-Fi network and probes alive supplicants in the WLAN. The attacker crafts ARP requests in the WLAN to identify alive supplicants [^1] Attackers can also identify alive supplicants by exploiting the DHCP mechanism, especially to circumvent the AP isolation mechanism enabled in Wi-Fi networks. Refer to Section 4.4 for more details.. By collecting the ARP replies, the attacker can obtain the $<MAC,IP>$ address pair of each alive supplicant which is also a potential victim client of our TCP hijacking attack. Then through analyzing the MAC address field of the captured wireless frames in the shared Wi-Fi channels, the attacker can filter the encrypted frames belonging to the victim client. If the Wi-Fi network provides multiple access channels, the attacker can scan all Wi-Fi channels to filter the victim’s frames. Once the victim’s frames are sniffed, the attacker gains a potent side channel to conduct the TCP hijacking attack.
@@
Third, the attacker infers the sequence number of the target TCP connection. The attacker pretends to be the victim supplicant and crafts TCP packets to the server. Those crafted TCP packets carry the guessed sequence numbers. If the guessed sequence number is less than the next sequence number to be received, the server will return a ACK packet carrying the SACK [^2] Selective acknowledgment (SACK) is an option in TCP that allows a receiver to acknowledge non-contiguous blocks of data received from the sender [24]. option in the TCP header to the supplicant. The SACK option in the TCP header will consume extra bits within the wireless frame. In contrast, if the attacker specifies a sequence number greater than the next sequence number, the return ACK packet from the server will not carry the SACK option. This subtle difference (i.e., the variation in frame size) can be observed by the attacker to infer the correct sequence number.
@@
Finally, we have responsibly reported this vulnerability to the Wi-Fi Alliance. Currently, we are discussing the mitigation measures with the Wi-Fi Alliance. The root cause of this vulnerability lies in the fixed size of Wi-Fi frames at the link layer, which inadvertently creates a reliable side channel and leaks information about TCP connections. As a result, we propose two possible countermeasures: (i) Modifying the 802.11 standards and dynamically padding the encrypted frames to prevent information leakage. (ii) Revisiting the TCP specifications so that the server responds consistently to different conditions.
@@ 2 Background
2.1 Frame Format and Security Mechanisms in Wi-Fi Network
@@ 2.2 Challenge ACK Mechanism in TCP
Our attack leverages the trigger conditions of the challenge ACK mechanism in two distinct ways. Firstly, when a receiver detects an incoming SYN packet within an established TCP connection, regardless of the sequence number, it responds by sending an ACK (referred to as the challenge ACK) to the remote peer. This ACK serves as a challenge for the remote peer to confirm the loss of the previous connection and the initiation of a new connection. Only the legitimate peer will receive this ACK and respond with a RST segment containing the correct sequence number, derived from the ACK field of the challenge ACK packet, in the event of connection loss. Consequently, a spoofed SYN packet will generate an additional ACK, which will be disregarded by the peer as a duplicate ACK and will have no impact on the established connection. We will demonstrate how this challenging condition can be exploited to detect a victim TCP connection in Section 4.2.
Secondly, the receiver employs a verification process for the acknowledgment number of each TCP segment to prevent blind data injection attacks. Acceptance of an acknowledgment number for any data segment is contingent upon its falling within the range of ($SND.UNA-SND.WND,SND.NXT$), where $SND.UNA$ represents the sequence number of the first unacknowledged octet, $SND.WND$ denotes the maximum window size observed by the receiver from the sender, and $SND.NXT$ is the next sequence number to be sent. If the acknowledgment number of the segment ($SEG.ACK$) is in the range ($SND.UNA-(2^{31}-1),SND.UNA-SND.WND$), the receiver responds with a challenge ACK. If the $SEG.ACK$ is greater than $SND.NXT$, the receiver silently discards this TCP segment. That can be exploited by attackers to infer the acceptable acknowledgment number, as described in Section 4.3.3.
Table 1: TCP packet size analysis with IPv4.
@@ 2.3 TCP Options
Timestamp Option. The TCP timestamp option is defined in RFC 1323 [12]. It is widely used in modern operating systems and various studies [44, 27, 31]. The timestamp option field spans a size of 10 octets, encompassing the timestamp value and timestamp echo reply fields. In practice, the timestamp option is typically padded with two extra bytes to maintain alignment of the TCP header on a 32-bit boundary. In a TCP connection with timestamp functionality enabled, the ACK packet includes a timestamp value indicating its transmission time. This timestamp can be utilized by the sender to calculate round-trip time and estimate the current network state. However, RST packets, which are employed for connection termination and lack TCP header options like the timestamp option, possess different sizes compared to ACK packets. This disparity in size between RST and ACK packets is illustrated in Table 1. In this paper, we will show that the different size of the ACK packet and the RST packet can be used to infer the source port number of a target TCP connection.
Selective Acknowledgment Option. The TCP selective acknowledgment (SACK) is specified in RFC 2018 [24]. It is an optional feature that is typically enabled by default in the majority of TCP implementations. The SACK option is particularly recommended for networks experiencing frequent packet loss or packet reordering. Its utilization can significantly improve the performance and reliability of TCP connections in such environments. When a receiver detects a TCP segment with a sequence number that has already been acknowledged as outdated, it responds by sending a SACK-ACK to notify the sender. The SACK option will enable the sender to selectively retransmit lost packets based on the information provided by the receiver. However, if the sequence number of the received TCP segment has not yet been acknowledged, the receiver will reply with an ACK packet, which may have a different frame size (as shown in Table 1) or may not respond at all, depending on the acknowledgment number of the segment. In this paper, we will show that attackers in Wi-Fi networks can differentiate between these situations and thus infer the sequence number of a target TCP connection by analyzing the size of encrypted wireless frames.
3 Threat Model
@@
4 TCP Hijacking with Encrypted Frame
Our attack exploits two key aspects. Firstly, the TCP stack exhibits inconsistent responses during packet verification. Depending on the validity of the received packet, the TCP receiver generates four different responses: no response packet, a RST, an ACK, and a SACK-ACK. Due to the presence of TCP options, these responses can be distinguished based on their packet sizes. Secondly, the frames within the Wi-Fi network are observable, and the frame sizes of these responses are consistently fixed (see Table 1). These characteristics create a significant side channel. An attacker can leverage this side channel to detect and hijack the victim’s TCP connection.
4.1 Attack Overview
@@
Step 2: Detecting TCP Connections. After detecting potential victims alive in the WLAN, the attacker impersonates the victim supplicant and sends forged SYN/ACK packets to the server. At the same time, the attacker monitors the victim’s encrypted frames in the Wi-Fi channel. By analyzing the encrypted frame size, the attacker can determine if a TCP connection exists between the victim and the server.
@@
4.2 Identifying victim and Detecting TCP Connections
@@
Detecting TCP Connections. With intercepting and analyzing the victim’s encrypted frames, the attacker can identify the victim’s TCP connections. A TCP connection is recognized by four elements, i.e., [client IP address, client port number, server IP address, server port number]. Typically, the server IP address and server port number are publicly known to the attacker [61, 14, 69]. The attacker needs to infer the client’s IP address and port number. In our attack, the client IP is obtained via ARP response. Thus, the only remaining element to determine is the client port number.
@@
4.3 Inferring Sequence and Acknowledgment Numbers
@@ 4.3.1 Verifying TCP Segment
4.3.2 Inferring the Exact Sequence Number
@@
To infer the exact sequence number on the server side, the attacker impersonates the client (i.e., a victim supplicant in WLAN) and sends forged TCP packets containing data to the server. These packets carry the guessed sequence numbers and a random acknowledgment number. The sequence number space is $2^{32}$ (i.e., 4G), and the server exhibits four distinct responses corresponding to different sequence numbers, as illustrated in Figure 4. (i) If the guessed sequence number falls within the range of ($RCV.NXT-2G,RCV.NXT-1$), the server returns a SACK-ACK response with an encrypted frame size of 80 bytes. (ii) If the guessed sequence number exceeds the upper boundary of the acceptable window (i.e., $RCV.NXT+RCV.WND$), the server sends an ACK response consisting of a 68-byte encrypted frame to the client. (iii) If the guessed sequence number is deemed acceptable but the random acknowledgment number ($SEQ.ACK$) is invalid (i.e., $SEQ.ACK>SND.NXT$), the server silently discards the packet. (iv) If the guessed sequence number is deemed acceptable and the random acknowledgment number falls within the challenge window, the server responds with a challenge ACK in compliance with RFC 5961 [59].
@@
4.3.3 Inferring an Acceptable Acknowledgment Number
@@
The challenge window for TCP segment acknowledgment number is defined in RFC 5961 [59] (see Section 2.2). As outlined in RFC 5961, the acknowledgment number space can be divided into three distinct cases, as shown in Figure 5. (i) The acknowledgment number falls within the challenge window, defined as ($SND.UNA-2G,SND.UNA-SND.WND$). (ii) The acknowledgment number resides within the acceptable ACK window, encompassing ($SND.UNA-SND.WND,SND.NXT$). (iii) Invalid acknowledgment numbers are those that exceed $SND.NXT$, denoted as $SEG.ACK>SND.NXT$.
@@ 4.4 Practical Considerations
Shifting Receive Window. When the victim’s TCP connection carries on ongoing traffic, the acceptable sequence and acknowledgment windows will shift during the attack. Fortunately, the attack can proceed as long as the inferred sequence number and acknowledgment number fall within the sliding window. The attacker can repeatedly infer the sequence number and acknowledgment number. Even if the receive window slides quickly enough to thwart the attacker’s inference, the attacker can opt to target the other end of the TCP connection. In this scenario, the receive window adjusts at a slower pace, as demonstrated in prior research [22, 14].
5 Case Study Attacks
In this section, we demonstrate two cases, i.e., SSH DoS and web manipulation, to illustrate how TCP connections can be hijacked by exploiting the encrypted frame size in Wi-Fi networks. In summary, an off-path attacker can reset an SSH service within 19 seconds and inject malicious data into a HTTP web page [^4] HTTPS can prevent attackers from injecting malicious data. However, reports on HTTPS adoption [67] indicate that there are 15% of websites still based on HTTP as of April 2024. within 28 seconds.
5.1 TCP DoS Attack
@@
Experimental Setup. This case involves three hosts: an SSH server (a rented VPS) running OpenSSH 8.4 and OpenSSL 1.1.1, a victim client (our laptop) running MacOS, and an attacker equipped with Kali 12.04 and multiple wireless network interface cards. The victim client is a supplicant in our Wi-Fi network and connects to the remote SSH server. The client will send commands to the server intermittently. Note that although the attacker and the victim supplicant are in the same Wi-Fi network, the attacker does not know the session key between the victim supplicant and the AP. The attacker attempts to terminate the connection by impersonating the victim supplicant and sending forged RST packets to the server. Taking into account the potential impact of Linux kernel versions, we strategically deploy servers with a variety of Linux kernel versions. Detailed configuration information for these servers is provided in Table 2.
Attack Procedure. In this attack, the off-path attacker needs to infer the 4-tuple of [client IP address, client port number, server IP address, server port number] and the exact sequence number of the target TCP connection. The server’s IP address and port number are publicly known to the attacker [61, 14, 69], thus it only needs to identify the other three remaining elements to proceed with the attack. The attacker first probes the victim client’s IP address and MAC address. Second, the attacker exploits the TCP header options to determine the client port number and infer the exact sequence number, as outlined in Section 4. Third, a crafted RST packet carrying the inferred value is issued to the server, and the server will be tricked into terminating the current SSH connection with the victim client.
Table 2: Experimental results of SSH connection reset.
@@
Results Evaluation. Table 2 displays the outcomes of our experiments, revealing that our attack is effective for different Linux versions. In particular, our attack exhibits an average bandwidth consumption of 75.76 KB/s, while maintaining an average execution time of 18.78 seconds. The empirical time cost distribution is shown in Figure 7. Our attack achieves a success rate of 84% on average. For the unsuccessful attempts, the primary cause is wireless interference, leading to the attacker missing crucial encrypted frames belonging to the victim. These frames contain the server’s responses to the probe packets. We will discuss wireless interference in depth in Section 7.
5.2 TCP Manipulation Attack

### v5

@@ Abstract
In this paper, we unveil a fundamental side channel in Wi-Fi networks, specifically the observable frame size, which can be exploited by attackers to conduct TCP hijacking attacks. Despite the various security mechanisms (e.g., WEP and WPA2/WPA3) implemented to safeguard Wi-Fi networks, our study reveals that an off-path attacker can still extract sufficient information from the frame size side channel to hijack the victim’s TCP connection. Our side channel attack is based on two significant findings: (i) response packets (e.g., ACK and RST) generated by TCP receivers vary in size, and (ii) the encrypted frames containing these response packets have consistent and distinguishable sizes. By observing the size of the victim’s encrypted frames, the attacker can detect and hijack the victim’s TCP connections. We validate the effectiveness of this side channel attack through two case studies, i.e., SSH DoS and web traffic manipulation. Precisely, our attack can terminate the victim’s SSH session in 19 seconds and inject malicious data into the victim’s web traffic within 28 seconds. Furthermore, we conduct extensive measurements to evaluate the impact of our attack on real-world Wi-Fi networks. We test 30 popular wireless routers from 9 well-known vendors, and none of these routers can protect victims from our attack. Besides, we implement our attack in 80 real-world Wi-Fi networks and successfully hijack the victim’s TCP connections in 75 (93.75%) evaluated Wi-Fi networks. We have responsibly disclosed the vulnerability to the Wi-Fi Alliance and proposed several mitigation strategies to address this issue.
publicationid: pubid: Network and Distributed System Security (NDSS) Symposium 2025 23 - 28 February 2025, San Diego, CA, USA ISBN 979-8-9894372-8-3 https://dx.doi.org/10.14722/ndss.2025.23305 www.ndss-symposium.org
I Introduction
@@
Our attack consists of four steps. First, the attacker accesses a public Wi-Fi network and probes alive supplicants in the WLAN. The attacker crafts ARP requests in the WLAN to identify alive supplicants [^1] Attackers can also identify alive supplicants by exploiting the DHCP mechanism, especially to circumvent the AP isolation mechanism enabled in Wi-Fi networks. Refer to Section IV-D for more details.. By collecting the ARP replies, the attacker can obtain the $<MAC,IP>$ address pair of each alive supplicant which is also a potential victim client of our TCP hijacking attack. Then through analyzing the MAC address field of the captured wireless frames in the shared Wi-Fi channels, the attacker can filter the encrypted frames belonging to the victim client. If the Wi-Fi network provides multiple access channels, the attacker can scan all Wi-Fi channels to filter the victim’s frames. Once the victim’s frames are sniffed, the attacker gains a potent side channel to conduct the TCP hijacking attack.
@@
Third, the attacker infers the sequence number of the target TCP connection. The attacker pretends to be the victim supplicant and crafts TCP packets to the server. Those crafted TCP packets carry the guessed sequence numbers. If the guessed sequence number is less than the next sequence number to be received, the server will return a ACK packet carrying the SACK [^2] Selective acknowledgment (SACK) is an option in TCP that allows a receiver to acknowledge non-contiguous blocks of data received from the sender. In this paper, we focus on exploiting the duplicate SACK option specified in RFC 2883 [48]. option in the TCP header to the supplicant. The SACK option in the TCP header will consume extra bits within the wireless frame. In contrast, if the attacker specifies a sequence number greater than the next sequence number, the return ACK packet from the server will not carry the SACK option. This subtle difference (i.e., the variation in frame size) can be observed by the attacker to infer the correct sequence number.
@@
Finally, we have responsibly reported this vulnerability to the Wi-Fi Alliance and they have acknowledged the issue. Currently, we are discussing the mitigation measures with the Wi-Fi Alliance. The root cause of this vulnerability lies in the fixed size of Wi-Fi frames at the link layer, which inadvertently creates a reliable side channel and leaks information about TCP connections. As a result, we propose two possible countermeasures: (i) Modifying the 802.11 standards and dynamically padding the encrypted frames to prevent information leakage. (ii) Revisiting the TCP specifications so that the server responds consistently to different conditions. The proof-of-concept (PoC) code for our attack is available at https://github.com/Internet-Architecture-and-Security/Packet-Size-Side-Channel-Attack.
@@ II Background
II-A Frame Format and Security Mechanisms in Wi-Fi Network
@@ II-B Challenge ACK Mechanism in TCP
Our attack leverages the trigger conditions of the challenge ACK mechanism in two distinct ways. Firstly, when a receiver detects an incoming SYN packet within an established TCP connection, regardless of the sequence number, it responds by sending an ACK (referred to as the challenge ACK) to the remote peer. This ACK serves as a challenge for the remote peer to confirm the loss of the previous connection and the initiation of a new connection. Only the legitimate peer will receive this ACK and respond with a RST segment containing the correct sequence number, derived from the ACK field of the challenge ACK packet, in the event of connection loss. Consequently, a spoofed SYN packet will generate an additional ACK, which will be disregarded by the peer as a duplicate ACK and will have no impact on the established connection. We will demonstrate how this challenging condition can be exploited to detect a victim TCP connection in Section IV-B.
Secondly, the receiver employs a verification process for the acknowledgment number of each TCP segment to prevent blind data injection attacks. Acceptance of an acknowledgment number for any data segment is contingent upon its falling within the range of ($SND.UNA-SND.WND,SND.NXT$), where $SND.UNA$ represents the sequence number of the first unacknowledged octet, $SND.WND$ denotes the maximum window size observed by the receiver from the sender, and $SND.NXT$ is the next sequence number to be sent, as illustrated for case (ii) in Figure 5. If the acknowledgment number of the segment ($SEG.ACK$) is in the range ($SND.UNA-(2^{31}-1),SND.UNA-SND.WND$), the receiver responds with a challenge ACK (see case (i) in Figure 5). If the $SEG.ACK$ is greater than $SND.NXT$, the receiver silently discards this TCP segment. That can be exploited by attackers to infer the acceptable acknowledgment number, as described in Section IV-C3.
TABLE I: TCP packet size analysis with IPv4.
@@ II-C TCP Options
Timestamp Option. The TCP timestamp option is defined in RFC 1323 [13]. It is widely used in modern operating systems and various studies [46, 28, 32]. The timestamp option field spans a size of 10 octets, encompassing the timestamp value and timestamp echo reply fields. In practice, the timestamp option is typically padded with two extra bytes to maintain alignment of the TCP header on a 32-bit boundary. In a TCP connection with timestamp functionality enabled, the ACK packet includes a timestamp value indicating its transmission time. This timestamp can be utilized by the sender to calculate round-trip time and estimate the current network state. However, RST packets, which are employed for connection termination and lack TCP header options like the timestamp option, possess different sizes compared to ACK packets. This disparity in size between RST and ACK packets is illustrated in Table I. In this paper, we will show that the different size of the ACK packet and the RST packet can be used to infer the source port number of a target TCP connection.
Selective Acknowledgment Option. The TCP selective acknowledgment (SACK) option is specified in RFC 2018 [25] and extended in RFC 2883 [48]. It is an optional feature that is typically enabled by default in the majority of TCP implementations. The SACK option is particularly recommended for networks experiencing frequent packet loss or packet reordering. Its utilization can significantly improve the performance and reliability of TCP connections in such environments. In this paper, we mainly discuss the extended SACK option (also known as duplicate SACK) specified in RFC 2883. This extension to the SACK option allows the TCP sender to infer the order of packets received at the receiver, allowing the sender to infer when it has unnecessarily retransmitted a packet. When a receiver detects a TCP segment with a sequence number that has already been acknowledged as outdated, it responds by sending a SACK-ACK to notify the sender. The sender could then use this information for more robust operations. However, if the sequence number of the received TCP segment has not yet been acknowledged, the receiver will reply with an ACK packet, which may have a different frame size (as shown in Table I) or may not respond at all, depending on the acknowledgment number of the segment. In this paper, we will show that attackers in Wi-Fi networks can differentiate between these situations and thus infer the sequence number of a target TCP connection by analyzing the size of encrypted wireless frames.
III Threat Model
@@
IV TCP Hijacking with Encrypted Frame
Our attack exploits two key aspects. Firstly, the TCP stack exhibits inconsistent responses during packet verification. Depending on the validity of the received packet, the TCP receiver generates four different responses: no response packet, a RST, an ACK, and a SACK-ACK. Due to the presence of TCP options, these responses can be distinguished based on their packet sizes. Secondly, the frames within the Wi-Fi network are observable, and the frame sizes of these responses are consistently fixed (see Table I). These characteristics create a significant side channel. An attacker can leverage this side channel to detect and hijack the victim’s TCP connection.
IV-A Attack Overview
@@
Step 2: Detecting TCP Connections. After detecting potential victims alive in the WLAN, the attacker impersonates the victim supplicant444“Impersonating the victim supplicant” refers to the attacker specifying the source IP address of crafted packets as the victim’s IP address. This works in WLANs where the AP does not check the IP addresses. and sends forged SYN/ACK packets to the server. At the same time, the attacker monitors the victim’s encrypted frames in the Wi-Fi channel. By analyzing the encrypted frame size, the attacker can determine if a TCP connection exists between the victim and the server.
@@
IV-B Identifying Victim and Detecting TCP Connections
@@
Detecting TCP Connections. With intercepting and analyzing the victim’s encrypted frames, the attacker can identify the victim’s TCP connections. Typically, the attacker focus on detecting TCP connections between the victim and popular servers [65, 15, 73], such as servers of famous websites. A TCP connection is recognized by four elements, i.e., [client IP address, client port number, server IP address, server port number]. The attacker can probe the server’s IP address (e.g., using “dig example.com”) and access the server to determine the server’s port number. Subsequently, the attacker needs to infer the client’s IP address and port number. In our attack, the client IP is obtained via ARP response. Thus, the last remaining element to determine is the client port number.
@@
IV-C Inferring Sequence and Acknowledgment Numbers
@@ IV-C1 Verifying TCP Segment
IV-C2 Inferring the Exact Sequence Number
@@
To infer the exact sequence number on the server side, the attacker impersonates the client (i.e., a victim supplicant in WLAN) and sends forged TCP packets containing data to the server. These packets carry the guessed sequence numbers and a random acknowledgment number. The sequence number space is $2^{32}$ (i.e., 4G), and the server exhibits four distinct responses corresponding to different sequence numbers [^5] The forged TCP packets with random acknowledgment numbers only elicit the server’s response with the duplicate SACK option., as illustrated in Figure 4. (i) If the guessed sequence number falls within the range of ($RCV.NXT-2G,RCV.NXT-1$), the server returns a SACK-ACK response with an encrypted frame size of 80 bytes. (ii) If the guessed sequence number exceeds the upper boundary of the acceptable window (i.e., $RCV.NXT+RCV.WND$), the server sends an ACK response consisting of a 68-byte encrypted frame to the client. (iii) If the guessed sequence number is deemed acceptable but the random acknowledgment number ($SEQ.ACK$) is invalid (i.e., $SEQ.ACK>SND.NXT$), the server silently discards the packet. (iv) If the guessed sequence number is deemed acceptable and the random acknowledgment number falls within the challenge window, the server responds with a challenge ACK in compliance with RFC 5961 [63].
@@
IV-C3 Inferring an Acceptable Acknowledgment Number
@@
The challenge window for TCP segment acknowledgment number is defined in RFC 5961 [63] (see Section II-B). As outlined in RFC 5961, the acknowledgment number space can be divided into three distinct cases, as shown in Figure 5. (i) The acknowledgment number falls within the challenge window, defined as ($SND.UNA-2G,SND.UNA-SND.WND$). (ii) The acknowledgment number resides within the acceptable ACK window, encompassing ($SND.UNA-SND.WND,SND.NXT$). (iii) Invalid acknowledgment numbers are those that exceed $SND.NXT$, denoted as $SEG.ACK>SND.NXT$.
@@ IV-D Practical Considerations
Shifting Receive Window. When the victim’s TCP connection carries on ongoing traffic, the acceptable sequence and acknowledgment windows will shift during the attack. Fortunately, the attack can proceed as long as the inferred sequence number and acknowledgment number fall within the sliding window. The attacker can repeatedly infer the sequence number and acknowledgment number. Even if the receive window slides quickly enough to thwart the attacker’s inference, the attacker can opt to target the other end of the TCP connection. In typical high-traffic scenarios like file downloading, the server-side sequence number triggers fast client-side acknowledgment, while the client-side sequence number grows slower. Attackers can infer the client-side sequence number and conduct brute-force attacks by sending multiple spoofed packets with acknowledgment numbers at different intervals, exploiting the large accepted window for acknowledgment number [43] at this stage.
V Case Study Attacks
In this section, we demonstrate two cases, i.e., SSH DoS and web manipulation, to illustrate how TCP connections can be hijacked by exploiting the encrypted frame size in Wi-Fi networks. In summary, an off-path attacker can reset an SSH service within 19 seconds and inject malicious data into a HTTP web page [^6] HTTPS can prevent attackers from injecting malicious data. However, reports on HTTPS adoption [71] indicate that there are 15% of websites still based on HTTP as of April 2024. Additionally, our measurements of the top 1 million websites based on Tranco [47] show that about 10% of them based on HTTP. within 28 seconds.
V-A TCP DoS Attack
@@
Experimental Setup. This case involves three hosts: an SSH server (a rented VPS) running OpenSSH 8.4 and OpenSSL 1.1.1, a victim client (our laptop) running MacOS, and an attacker equipped with Kali 2023.1 and multiple wireless network interface cards. The victim client is a supplicant in our Wi-Fi network and connects to the remote SSH server. The client will send commands to the server intermittently. Note that although the attacker and the victim supplicant are in the same Wi-Fi network, the attacker does not know the session key between the victim supplicant and the AP. The attacker attempts to terminate the connection by impersonating the victim supplicant and sending forged RST packets to the server. Taking into account the potential impact of Linux kernel versions, we strategically deploy servers with a variety of Linux kernel versions. Detailed configuration information for these servers is provided in Table II.
Attack Procedure. In this attack, the off-path attacker needs to infer the 4-tuple of [client IP address, client port number, server IP address, server port number] and the exact sequence number of the target TCP connection. The server’s IP address and port number are publicly known to the attacker [65, 15, 73], thus it only needs to identify the other three remaining elements to proceed with the attack. The attacker first probes the victim client’s IP address and MAC address. Second, the attacker exploits the TCP header options to determine the client port number and infer the exact sequence number, as outlined in Section IV. Third, a crafted RST packet carrying the inferred value is issued to the server, and the server will be tricked into terminating the current SSH connection with the victim client.
TABLE II: Experimental results of SSH connection reset.
@@
Results Evaluation. Table II displays the outcomes of our experiments, revealing that our attack is effective for different Linux versions. In particular, our attack exhibits an average bandwidth consumption of 75.76 KB/s, while maintaining an average execution time of 18.78 seconds. The empirical time cost distribution is shown in Figure 7. Our attack achieves a success rate of 84% on average. For the unsuccessful attempts, the primary cause is wireless interference, leading to the attacker missing crucial encrypted frames belonging to the victim. These frames contain the server’s responses to the probe packets. We will discuss wireless interference in depth in Section VII.
V-B TCP Manipulation Attack